# Approaching Stable Quark Matter

### v1

@@ Appendix B Integral approximations
$I_{13}(\hat{m})$ $\simeq\frac{8}{3}(1-\ln 2)\hat{u}^{3}+0.318\hat{u}^{6}-0.137\hat{u}^{7}\leavevmode\nobreak\ ,$ (B.1)
$I_{14}(\hat{m})$ $\simeq 1.99\hat{v}-0.99\hat{v}^{2}+\ln\hat{v}\left(-0.27\hat{v}+0.26\hat{v}^{2}-0.119\hat{v}^{2}\ln\hat{v}\right)\leavevmode\nobreak\ ,$ (B.2)
@@
$+4\mu_{i}\mu_{j}(\mu_{i}^{2}+\mu_{j}^{2})\ln\frac{(\mu_{i}+\mu_{j})^{2}}{\mu_{i}\mu_{j}}-(\mu_{i}^{4}-\mu_{j}^{4})\ln\frac{\mu_{i}}{\mu_{j}}\Bigg{]}\leavevmode\nobreak\ ,$ (B.3)
$I_{16}(\hat{m},\vec{\hat{\mu}}^{2})$ $=-0.8564+2\ln\left(\frac{\vec{\hat{\mu}}^{4}+2\vec{\hat{\mu}}^{2}I_{14}+I_{14}I_{16h1}}{\vec{\hat{\mu}}^{4}+\vec{\hat{\mu}}^{2}I_{14}}\right)\leavevmode\nobreak\ ,$ (B.4)
$I_{16h1}(\hat{m})$ $\simeq 4.629\hat{v}-3.629\hat{v}^{2}+\ln\hat{v}\left(0.6064\hat{v}+2.021\hat{v}^{2}-0.566\hat{v}^{2}\ln\hat{v}\right)\leavevmode\nobreak\ ,$ (B.5)
@@
$+I_{17h2}\frac{\hat{\mu}_{i}}{\hat{\mu}_{i}+1}+I_{17h3}\frac{1}{\hat{\mu}_{i}^{2}+1}+I_{17h4}\frac{\hat{\mu}_{i}}{3.38\sqrt{1-\hat{m}}+7.96(1-\hat{m})+\hat{\mu}_{i}^{3}}\leavevmode\nobreak\ ,$ (B.6)
$I_{17h1}(\hat{m})$ $\simeq-7.5439\hat{v}+1.1422\hat{u}+6.639\hat{u}^{5}-8.8809\hat{u}^{3}\ln(1+\hat{m})$
$+\ln\hat{v}\left(-2.005\hat{v}+5.865\hat{v}^{2}\ln\hat{v}-10.622\hat{v}^{3}\ln^{2}\hat{v}\right)\leavevmode\nobreak\ ,$ (B.7)
$I_{17h2}(\hat{m})$ $\simeq-8.57\hat{v}+3.09\hat{u}+9.654\hat{u}^{2}+10.86\hat{u}^{3}-12.445\hat{u}^{4}$
$+\hat{u}\ln\hat{v}(0.374+8.73\hat{u}^{2})-\frac{7}{3}I_{14}-I_{17h1}\leavevmode\nobreak\ ,$ (B.8)
$I_{17h3}(\hat{m})$ $\simeq 4.804\hat{v}-2.224\hat{v}^{2}+\ln\hat{v}(-0.411\hat{v}+0.056\hat{v}^{2}-0.0718\hat{v}^{2}\ln\hat{v})-\frac{7}{3}I_{14}-I_{17h1}\leavevmode\nobreak\ ,$ (B.9)
$I_{17h4}(\hat{m})$ $\simeq-8.085\hat{v}^{3}+6.646\hat{v}^{5}-0.586\hat{v}^{12}+0.3465\hat{m}\hat{u}$
$+1.359\hat{u}^{2}-3.77\hat{u}^{3}+5.395\hat{u}^{4}-0.9525\hat{u}^{5}\leavevmode\nobreak\ ,$ (B.10)
$I_{18}(\hat{m})$ $\simeq 2.3\hat{v}-1.3\hat{v}^{2}+\ln\hat{v}\left(-0.3477\hat{v}+0.6484\hat{v}^{2}-0.241\hat{v}^{2}\ln\hat{v}\right)\leavevmode\nobreak\ ,$ (B.11)
$I_{19}(\hat{m},\vec{\hat{\mu}}^{2})$ $=I_{19h1}+I_{14}\left[2\ln\vec{\hat{\mu}}^{2}+2\ln\left(\frac{I_{14}\vec{\hat{\mu}}^{4}+2I_{13}\vec{\hat{\mu}}^{2}+I_{14}I_{19h2}}{I_{14}\vec{\hat{\mu}}^{4}+I_{13}\vec{\hat{\mu}}^{2}}\right)\right]\leavevmode\nobreak\ ,$ (B.12)
$I_{19h1}(\hat{m})$ $\simeq-3.6643\hat{v}+2.808\hat{v}^{2}+\ln\hat{v}(-0.7305\hat{v}-1.22\hat{v}^{2})$
$+\ln^{2}\hat{v}(-0.0437\hat{v}+0.312\hat{v}^{2}-0.0237\hat{v}^{2}\ln\hat{v})\leavevmode\nobreak\ ,$ (B.13)
$I_{19h2}(\hat{m})$ $\simeq 0.104\hat{v}+0.464\hat{v}^{2}+0.432\hat{u}^{5}+\ln\hat{v}(-0.04\hat{v}-0.99\hat{v}^{2}+0.455\hat{v}^{2}\ln\hat{v})\leavevmode\nobreak\ ,$ (B.14)
$I_{20}(\hat{m})$ $\simeq 3.84\hat{v}-7.94\hat{v}^{2}+8.06\hat{u}^{2}+\ln\hat{v}(3.81\hat{v}+9.7\hat{v}^{2}-7.42\hat{v}^{2}\ln\hat{v})\leavevmode\nobreak\ ,$ (B.15)
$I_{21}(\hat{m})$ $\simeq-0.11\hat{v}+1.11\hat{v}^{2}+\ln\hat{v}(-0.059\hat{v}-2.07\hat{v}^{2}+0.4\hat{v}^{2}\ln\hat{v})\leavevmode\nobreak\ ,$ (B.16)
$I_{22}(\hat{m},\vec{\hat{\mu}}^{2})$ $=I_{22h1}+I_{13}\left[2\ln\vec{\hat{\mu}}^{2}+2\ln\left(\frac{\vec{\hat{\mu}}^{4}+\vec{\hat{\mu}}^{2}I_{22h2}+I_{22h3}}{\vec{\hat{\mu}}^{4}+0.5\vec{\hat{\mu}}^{2}I_{22h2}}\right)\right]\leavevmode\nobreak\ ,$ (B.17)
$I_{22h1}(\hat{m})$ $\simeq-0.8255\hat{u}^{2}-0.03084\hat{u}^{7}+0.00161\hat{v}\ln\hat{v}\leavevmode\nobreak\ ,$ (B.18)
$I_{22h2}(\hat{m})$ $\simeq 1.313\hat{u}+0.434\hat{u}^{2}+0.253\hat{u}^{5}\leavevmode\nobreak\ ,$ (B.19)
$I_{22h3}(\hat{m})$ $\simeq 1.844\hat{v}-0.844\hat{v}^{2}+\ln\hat{v}(0.194\hat{v}-0.392\hat{v}^{2}-0.704\hat{v}^{2}\ln\hat{v})\leavevmode\nobreak\ .$ (B.20)
@@ Appendix C Formulas of color superconductivity
$p_{\rm CS}=d\frac{\mu^{2}}{4\pi^{2}}\Delta^{2}\left[1+\frac{2\pi}{3^{3/2}\sqrt{c_{R}}}\left(\frac{\alpha_{s}}{\pi}\right)^{1/2}+\mathcal{O}(\alpha_{s})\right]\leavevmode\nobreak\ ,$ (C.1)
@@
$\ln\left(\frac{\Delta}{\mu}\right)=-\frac{\sqrt{3}\,\pi}{2\,\sqrt{c_{R}}}\left(\frac{\alpha_{s}}{\pi}\right)^{-1/2}-\frac{5}{2}\ln\left(N_{f}\frac{\alpha_{s}}{\pi}\right)+\ln\frac{2^{13/2}}{\pi}-\frac{\pi^{2}+4}{12\,c_{R}}-\zeta+\mathcal{O}(\alpha_{s}^{1/2})\leavevmode\nobreak\ ,$ (C.2)
where $\zeta=\frac{1}{3}\ln 2\,(0)$, $c_{R}=2/3\,(4/3)$, and $d=12\,(6)$ for the CFL (color singlet $\bar{q}q$) phases in baryon-dense (isospin-dense) matter, respectively. Because of the different $c_{R}$ factors, $p_{\rm CS}$ for the CFL channel is much more suppressed compared to the pQCD contributions, while $p_{\rm CS}$ for the color singlet $\bar{q}q$ channel plays a major role in the analysis of the isospin-dense matter (see Section 4).
@@ Appendix D Additional details of the Lattice QCD data analysis
We also present the results of the same analysis done in Sections 4 and 5 for $\mu_{\rm start}=800$ MeV in Figures 10 and 11. The $1\sigma$ and $90\%$ CL best-fit parameters are given by $(X,B^{1/4})=(1.30,0\leavevmode\nobreak\ {\rm MeV})$ and $(X,B^{1/4})=(1.58,0\leavevmode\nobreak\ {\rm MeV})$, while we also present the benchmarks of $(X,B^{1/4})=(1.30,176\leavevmode\nobreak\ {\rm MeV})$ (red dashed line) which sits on the $90\%$ CL contour around the $k=30$ best-fit point, $(X,B^{1/4})=(1.30,200\leavevmode\nobreak\ {\rm MeV})$ (red dotted line), and $(1.30,250\leavevmode\nobreak\ {\rm MeV})$ (red dot-dashed line).
Figure 10: The $c_{s}^{2}$, $p/p_{\rm id}$, and $\epsilon/\epsilon_{\rm id}$ distributions of the best-fit points (solid lines) with $\mu_{\rm start}=800$ MeV and $k=30,35$, whose benchmark parameters are $(X,B^{1/4})=(1.30,0\leavevmode\nobreak\ {\rm MeV})$ and $(X,B^{1/4})=(1.58,0\leavevmode\nobreak\ {\rm MeV})$. We also show the distributions of the benchmarks $(X,B^{1/4})=(1.30,176\leavevmode\nobreak\ {\rm MeV})$ (red dashed line) which sits on the $90\%$ CL contour around the $k=30$ best-fit point, $(X,B^{1/4})=(1.30,200\leavevmode\nobreak\ {\rm MeV})$ (red dotted line), and $(1.30,250\leavevmode\nobreak\ {\rm MeV})$ (red dot-dashed line). The red lines completely overlap with one another for $c_{s}^{2}$ since it does not depend on $B$.
Figure 11: The $90\%$ CL quark matter stability constraint for $\mu_{\rm start}=800$ MeV and $k=30,35$ (solid lines) as well as the corresponding future projection with $N_{\rm scale}=\sqrt{10}$ (dot-dashed lines).
@@ Appendix E Thermodynamic bounds on the QCD EOS
    $\partial_{\mu}^{2}\Omega(\mu)=-\partial_{\mu}^{2}p(\mu)=-\partial_{\mu}n(\mu)\leq 0\to\partial_{\mu}n(\mu)\geq 0\leavevmode\nobreak\ .$ (E.1)
@@
[…]
@@
[…]
@@
$n_{\rm min}(\mu)\leq n(\mu)\leq n_{\rm max}(\mu)\leavevmode\nobreak\ ,$ (E.4)
[…]
@@
$\mu_{c}=\sqrt{\frac{\mu_{L}\,\mu_{H}\,(\mu_{H}\,n_{H}-\mu_{L}\,n_{L}-2\,\Delta p)}{\mu_{L}\,n_{H}-\mu_{H}\,n_{L}}}\leavevmode\nobreak\ .$ (E.7)
@@
$p_{\rm min}(\mu,n)\leq p(\mu,n)\leq p_{\rm max}(\mu,n)\leavevmode\nobreak\ ,$ (E.8)
$p_{\rm min}(\mu,n)$ $=p_{L}+\frac{\mu^{2}-\mu_{L}^{2}}{2\mu}n_{\rm min}(\mu)\leavevmode\nobreak\ ,$ (E.9)
[…]
@@
$n_{c}(\mu)=\frac{n_{\rm max}(\mu_{L})}{\mu_{L}}\mu=\frac{n_{\rm min}(\mu_{H})}{\mu_{H}}\mu\leavevmode\nobreak\ .$ (E.11)
@@
$p_{\rm min}(\mu)$ $=p_{\rm min}\big{(}\mu,n_{\rm min}(\mu)\big{)}\leavevmode\nobreak\ ,$ (E.12)
$p_{\rm max}(\mu)$ $=p_{\rm max}\big{(}\mu,n_{c}(\mu)\big{)}\leavevmode\nobreak\ .$ (E.13)
@@ Appendix F Quark matter EOS and ground state energy
In this section, we study the thermodynamic constraints on the quark matter EOS. While in principle one can choose $\mu_{L}=\mu_{0}>0$ for the low-density reference point when constraining the quark matter EOS, the uncertainty of pQCD calculations at that point is very high and thus unreliable (see Sections 2 and 3). A more conservative choice is to choose $\mu_{L}=0$ with $n_{L}=0$ and will lead to a less stringent but more reliable bound, which we use in this study. To specify $p_{L}$, we need additional inputs about the deconfined vacuum in terms of the bag parameter $B$, with which we can set $p_{L}=-B$. On the other hand, the high-density reference point can be obtained through pQCD calculations, and one should also shift $p_{{\rm pQCD},H}\to p_{H}=p_{{\rm pQCD},H}-B$ to include the nonperturbative corrections. We show one specific example of $p_{L}=-(150\leavevmode\nobreak\ {\rm MeV})^{4}$ and $(\mu_{H},n_{H},p_{H})$ $=$ $(1500\leavevmode\nobreak\ {\rm MeV},0.916\leavevmode\nobreak\ {\rm fm^{-3}},322.6\leavevmode\nobreak\ {\rm MeV/fm^{3}})$, one benchmark of the 2+1-flavor $X=1.6$ scenario, in Figure 12.
Figure 12: Thermodynamic bounds on (left) $n(\mu)$ and (right) $p(\mu)$ using the 2+1-flavor $X=1.6$ benchmark for the high-density reference point with $\mu_{H}=1500$ MeV and $B=(150\leavevmode\nobreak\ {\rm MeV})^{4}$. In the right panel, the green crosses are the intersections between the thermodynamics bounds and the $p=0$ line, which give the minimally and maximally allowed quark matter ground state energies per baryon ($E_{Q,{\rm min,max}}(B)=\mu_{Q,{\rm min,max}}(B)$), respectively.
@@
$\left.\frac{\partial E}{\partial\mu}\right|_{\rm\mu=\mu_{Q}}=-\frac{n^{2}-p(\mu_{Q})(\partial n/\partial\mu)_{\mu=\mu_{Q}}}{n^{2}}+1=0\Rightarrow p(\mu_{Q})=0\leavevmode\nobreak\ {\rm or}\leavevmode\nobreak\ \left(\frac{\partial n}{\partial\mu}\right)_{\mu=\mu_{Q}}=0\leavevmode\nobreak\ .$ (F.1)

### v3

@@ Appendix B Integral approximations
[…]
@@
$+4\mu_{i}\mu_{j}(\mu_{i}^{2}+\mu_{j}^{2})\ln\frac{(\mu_{i}+\mu_{j})^{2}}{\mu_{i}\mu_{j}}-(\mu_{i}^{4}-\mu_{j}^{4})\ln\frac{\mu_{i}}{\mu_{j}}\Bigg]\penalty 10000\ ,$ (B.3)
$I_{16}(\hat{m},\vec{\hat{\mu}}^{2})$ $=-0.8564+2\ln\left(\frac{\vec{\hat{\mu}}^{4}+2\vec{\hat{\mu}}^{2}I_{14}+I_{14}I_{16h1}}{\vec{\hat{\mu}}^{4}+\vec{\hat{\mu}}^{2}I_{14}}\right)\penalty 10000\ ,$ (B.4)
[…]
@@
$+I_{17h2}\frac{\hat{\mu}_{i}}{\hat{\mu}_{i}+1}+I_{17h3}\frac{1}{\hat{\mu}_{i}^{2}+1}+I_{17h4}\frac{\hat{\mu}_{i}}{3.38\sqrt{1-\hat{m}}+7.96(1-\hat{m})+\hat{\mu}_{i}^{3}}\penalty 10000\ ,$ (B.6)
[…]
$+\hat{u}\ln\hat{v}(0.374+8.73\hat{u}^{2})-\frac{7}{3}I_{14}-I_{17h1}\penalty 10000\ ,$ (B.8)
[…]
$+1.359\hat{u}^{2}-3.77\hat{u}^{3}+5.395\hat{u}^{4}-0.9525\hat{u}^{5}\penalty 10000\ ,$ (B.10)
[…]
$I_{19}(\hat{m},\vec{\hat{\mu}}^{2})$ $=I_{19h1}+I_{14}\left[2\ln\vec{\hat{\mu}}^{2}+2\ln\left(\frac{I_{14}\vec{\hat{\mu}}^{4}+2I_{13}\vec{\hat{\mu}}^{2}+I_{14}I_{19h2}}{I_{14}\vec{\hat{\mu}}^{4}+I_{13}\vec{\hat{\mu}}^{2}}\right)\right]\penalty 10000\ ,$ (B.12)
[…]
$I_{22}(\hat{m},\vec{\hat{\mu}}^{2})$ $=I_{22h1}+I_{13}\left[2\ln\vec{\hat{\mu}}^{2}+2\ln\left(\frac{\vec{\hat{\mu}}^{4}+\vec{\hat{\mu}}^{2}I_{22h2}+I_{22h3}}{\vec{\hat{\mu}}^{4}+0.5\vec{\hat{\mu}}^{2}I_{22h2}}\right)\right]\penalty 10000\ ,$ (B.17)
[…]
@@ Appendix C Formulas of color superconductivity
$p_{\rm CS}=d\frac{\mu^{2}}{4\pi^{2}}\Delta^{2}\left[1+\frac{2\pi}{3^{3/2}\sqrt{c_{R}}}\left(\frac{\alpha_{s}}{\pi}\right)^{1/2}+\mathcal{O}(\alpha_{s})\right]\penalty 10000\ ,$ (C.1)
@@
$\ln\left(\frac{\Delta}{\mu}\right)=-\frac{\sqrt{3}\,\pi}{2\,\sqrt{c_{R}}}\left(\frac{\alpha_{s}}{\pi}\right)^{-1/2}-\frac{5}{2}\ln\left(N_{f}\frac{\alpha_{s}}{\pi}\right)+\ln\frac{2^{13/2}}{\pi}-\frac{\pi^{2}+4}{12\,c_{R}}-\zeta+\mathcal{O}(\alpha_{s}^{1/2})\penalty 10000\ ,$ (C.2)
where $\zeta=\frac{1}{3}\ln 2\,(0)$, $c_{R}=2/3\,(4/3)$, and d=12(6)d=12\,(6) for the CFL (color singlet $\bar{q}q$) phases in baryon-dense (isospin-dense) matter, respectively. Because of the different $c_{R}$ factors, $p_{\rm CS}$ for the CFL channel is much more suppressed compared to the pQCD contributions, while $p_{\rm CS}$ for the color singlet $\bar{q}q$ channel plays a major role in the analysis of the isospin-dense matter (see Section 4).
@@ Appendix D Additional details of the Lattice QCD data analysis
We also present the results of the same analysis done in Sections 4 and 5 for $\mu_{\rm start}=800$ MeV in Figures 12 and 13. The $1\sigma$ and $90\%$ CL best-fit parameters are given by $(X,B^{1/4})=(1.30,0\penalty 10000\ {\rm MeV})$ and $(X,B^{1/4})=(1.58,0\penalty 10000\ {\rm MeV})$, while we also present the benchmarks of $(X,B^{1/4})=(1.30,176\penalty 10000\ {\rm MeV})$ (red dashed line) which sits on the $90\%$ CL contour around the $k=30$ best-fit point, $(X,B^{1/4})=(1.30,200\penalty 10000\ {\rm MeV})$ (red dotted line), and $(1.30,250\penalty 10000\ {\rm MeV})$ (red dot-dashed line).
Figure 12: The $c_{s}^{2}$, $p/p_{\rm id}$, and $\epsilon/\epsilon_{\rm id}$ distributions of the best-fit points (solid lines) with $\mu_{\rm start}=800$ MeV and $k=30,35$, whose benchmark parameters are $(X,B^{1/4})=(1.30,0\penalty 10000\ {\rm MeV})$ and $(X,B^{1/4})=(1.58,0\penalty 10000\ {\rm MeV})$. We also show the distributions of the benchmarks $(X,B^{1/4})=(1.30,176\penalty 10000\ {\rm MeV})$ (red dashed line) which sits on the $90\%$ CL contour around the $k=30$ best-fit point, $(X,B^{1/4})=(1.30,200\penalty 10000\ {\rm MeV})$ (red dotted line), and $(1.30,250\penalty 10000\ {\rm MeV})$ (red dot-dashed line). The red lines completely overlap with one another for $c_{s}^{2}$ since it does not depend on $B$.
Figure 13: The $90\%$ CL (left) 2+1-flavor and (right) 2-flavor quark matter stability constraint for $\mu_{\rm start}=800$ MeV and $k=30,35$ (solid lines) as well as the corresponding future projection with $N_{\rm scale}=\sqrt{10}$ (dot-dashed lines). The lower bounds from Eqs. (5.3) and (5.4) are also shown in the left and right panels, respectively.
@@ Appendix E Thermodynamic bounds on the QCD EOS
    $\partial_{\mu}^{2}\Omega(\mu)=-\partial_{\mu}^{2}p(\mu)=-\partial_{\mu}n(\mu)\leq 0\to\partial_{\mu}n(\mu)\geq 0\penalty 10000\ .$ (E.1)
@@
    $\text{Causality constraint:}\penalty 10000\ c_{s}^{-2}=\frac{\mu}{n}\frac{\partial n}{\partial\mu}\geq 1\penalty 10000\ .\hskip 62.59596pt$ (E.2)
@@
    $\text{Integral constraint:}\penalty 10000\ \int_{\mu_{L}}^{\mu_{H}}n(\mu)d\mu=p_{H}-p_{L}=\Delta p\penalty 10000\ .$ (E.3)
@@
$n_{\rm min}(\mu)\leq n(\mu)\leq n_{\rm max}(\mu)\penalty 10000\ ,$ (E.4)
$n_{\rm min}(\mu)$ $=\begin{cases}\begin{aligned} &\frac{n_{L}}{\mu_{L}}\mu&,\penalty 10000\ \quad\mu_{L}\leq\mu\leq\mu_{c}\\[4.30554pt] &\frac{\mu^{3}\,n_{H}-\mu\,\mu_{H}(\mu_{H}\,n_{H}-2\,\Delta p)}{(\mu^{2}-\mu_{L}^{2})\,\mu_{H}}&,\penalty 10000\ \quad\mu_{c}<\mu\leq\mu_{H}\end{aligned}\end{cases}\penalty 10000\ ,$ (E.5)
$n_{\rm max}(\mu)$ $=\begin{cases}\begin{aligned} &\frac{\mu^{3}\,n_{L}-\mu\,\mu_{L}(\mu_{L}\,n_{L}+2\,\Delta p)}{(\mu^{2}-\mu_{H}^{2})\,\mu_{L}}&,\penalty 10000\ \quad\mu_{L}\leq\mu<\mu_{c}\\[4.30554pt] &\frac{n_{H}}{\mu_{H}}\mu&,\penalty 10000\ \quad\mu_{c}\leq\mu\leq\mu_{H}\end{aligned}\end{cases}\penalty 10000\ ,$ (E.6)
@@
$\mu_{c}=\sqrt{\frac{\mu_{L}\,\mu_{H}\,(\mu_{H}\,n_{H}-\mu_{L}\,n_{L}-2\,\Delta p)}{\mu_{L}\,n_{H}-\mu_{H}\,n_{L}}}\penalty 10000\ .$ (E.7)
@@
$p_{\rm min}(\mu,n)\leq p(\mu,n)\leq p_{\rm max}(\mu,n)\penalty 10000\ ,$ (E.8)
$p_{\rm min}(\mu,n)$ $=p_{L}+\frac{\mu^{2}-\mu_{L}^{2}}{2\mu}n_{\rm min}(\mu)\penalty 10000\ ,$ (E.9)
$p_{\rm max}(\mu,n)$ $=\begin{cases}\begin{aligned} &p_{L}+\frac{\mu^{2}-\mu_{L}^{2}}{2\mu}n&,\penalty 10000\ \quad n\leq n_{c}(\mu)\\[4.30554pt] &p_{H}-\frac{\mu_{H}^{2}-\mu^{2}}{2\mu}n&,\penalty 10000\ \quad n>n_{c}(\mu)\end{aligned}\penalty 10000\ ,\end{cases}$ (E.10)
@@
$n_{c}(\mu)=\frac{n_{\rm max}(\mu_{L})}{\mu_{L}}\mu=\frac{n_{\rm min}(\mu_{H})}{\mu_{H}}\mu\penalty 10000\ .$ (E.11)
@@
$p_{\rm min}(\mu)$ $=p_{\rm min}\big(\mu,n_{\rm min}(\mu)\big)\penalty 10000\ ,$ (E.12)
$p_{\rm max}(\mu)$ $=p_{\rm max}\big(\mu,n_{c}(\mu)\big)\penalty 10000\ .$ (E.13)
@@ Appendix F Quark matter EOS and ground state energy
In this section, we study the thermodynamic constraints on the quark matter EOS. While in principle one can choose $\mu_{L}=\mu_{0}>0$ for the low-density reference point when constraining the quark matter EOS, the uncertainty of pQCD calculations at that point is very high and thus unreliable (see Sections 2 and 3). A more conservative choice is to choose $\mu_{L}=0$ with $n_{L}=0$ and will lead to a less stringent but more reliable bound, which we use in this study. To specify $p_{L}$, we need additional inputs about the deconfined vacuum in terms of the bag parameter $B$, with which we can set $p_{L}=-B$. On the other hand, the high-density reference point can be obtained through pQCD calculations, and one should also shift $p_{{\rm pQCD},H}\to p_{H}=p_{{\rm pQCD},H}-B$ to include the nonperturbative corrections. We show one specific example of $p_{L}=-(150\penalty 10000\ {\rm MeV})^{4}$ and $(\mu_{H},n_{H},p_{H})$ == $(1400\penalty 10000\ {\rm MeV},0.469\penalty 10000\ {\rm fm^{-3}},112.6\penalty 10000\ {\rm MeV/fm^{3}})$, one benchmark of the 2-flavor $X=1.6$ scenario, in Figure 14.
Figure 14: Thermodynamic bounds on (left) $n(\mu)$ and (right) $p(\mu)$ using the 2-flavor $X=1.6$ benchmark for the high-density reference point with $\mu_{H}=1400$ MeV and $B=(150\penalty 10000\ {\rm MeV})^{4}$. In the right panel, the green crosses are the intersections between the thermodynamics bounds and the $p=0$ line, which give the minimally and maximally allowed quark matter ground state energies per baryon ($E_{Q,{\rm min,max}}(B)=\mu_{Q,{\rm min,max}}(B)$), respectively.
@@
$\left.\frac{\partial E}{\partial\mu}\right|_{\rm\mu=\mu_{Q}}=-\frac{n^{2}-p(\mu_{Q})(\partial n/\partial\mu)_{\mu=\mu_{Q}}}{n^{2}}+1=0\Rightarrow p(\mu_{Q})=0\penalty 10000\ {\rm or}\penalty 10000\ \left(\frac{\partial n}{\partial\mu}\right)_{\mu=\mu_{Q}}=0\penalty 10000\ .$ (F.1)